40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7715     Entry time: Thu Nov 15 03:09:08 2012
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: No good progress, many options 
I didn't make any concrete progress today. AS and REFL cameras are in place, and Manasa has put ND 0.5 filters on both. I used a 
camera to look at the back of the Faraday, and aligned PRM such that it was retroreflecting, and then tried to align ITMY to have once 
fringes with the PRM at that place. I failed in this, since the AS beam on the AS table was starting to dall off the first mirror on the table. 
I then restored all the suspensions to where they were before I started touching them today. 

I moved ETMY face camera so that it is looking at the front of the black glass, but it's hard to tell where the beam is.  I was thinking 
about setting up a temporary camera to look at the back of ITMY to help guide PZT steering, but I haven't done this yet. 

Koji and I then talked about the several different options I have for references, and how many different knobs I  can turn. I'm sleeping 
on it for now, and hopefully I'll have more insight on what to do tomorrow. 
ELOG V3.1.3-